STAT 115 Let X be a continuous random variable having the CDF

Fx(x) = 1 - e^ (-e^x)

(1) Find the Probability Density Function (PDF) of Y=e^X.

(2) Let B have a uniform distribution over (0,1). Find a function G(b) and G(B) has the same distribution as X.
Answer #1

1)

P(Y < y)

= P(e^x < y)

= P(X < ln y )

= F(ln y)

= 1 - e^ (-e^(ln y))

= 1- e^(-y))

hence

f(y) = d/dy F(y) = e^(-y)

Y follow exponential distribution with mean = 1

Please give me a thumbs-up if this helps you out. Thank you! :)

By HomeworkLib answering guidelines, we have to answer only one question at a time